(The following payments are the full amount of the loss of a company. 27, 115, 126, 155, 163,243, 357, 506, 701, 877, 1193, 1340, 1884, 2598, 117743 a.) Empirically estimate the probability that a payment will be at least 730, when there is a deductible of 420

b.  Estimate the variance of the estimator in (a) above
Answer #1

a).

P(payment >= 730) = (no. of payment >= 730)/(total no. of payments)

= 6 / 15

= 0.4

b).

variance of p = [p*(1-p)/n]

= 0.4*(1-0.4)/15

= 0.4*0.6/15

= 0.016

variance = 0.016

--------------------------------------

DEAR STUDENT,

IF YOU HAVE ANY QUERY ASK ME IN THE COMMENT BOX,I AM HERE TO HELPS YOU.PLEASE GIVE ME POSITIVE RATINGS

*****************THANK YOU***************